OpenStudy (anonymous):

Need help to figure out an antiderivative. squareroot(4-x^2)

OpenStudy (anonymous):

\[\int\limits_{0}^{b}16\sqrt{4-x ^{2}}\]

OpenStudy (anonymous):

if forgot to add the 16

OpenStudy (anonymous):

we're working with volumes and we want to find one of the endpoint b that would give a volume of 19.654

OpenStudy (anonymous):

i figure that by using the fundamental theorem i could solve for that endpoint

OpenStudy (raden):

using trig-sub it will work first, let x=2sin(theta) what happend next ?

OpenStudy (anonymous):

we have \[16\sqrt{4-4\sin ^{2} \theta}\]

OpenStudy (raden):

simplify, can be to sqrt(4-4sin^2 theta) = sqrt(4(1-sin^2 theta)) = sqrt(4cos^2 theta) = 2cos(theta) note : sin^2 (theta) + cos^2 (theta) = 1

OpenStudy (raden):

and because x=2sin(theta) (differentiate both side) so, dx=2cos(theta) d(theta)

OpenStudy (anonymous):

\[\int\limits_{0}^{b}\frac{ 16 }{ \sqrt{4-x^2} }\]

OpenStudy (anonymous):

i think your question is like this

OpenStudy (raden):

but i dont see divide at first

OpenStudy (anonymous):

03453660 hmm, no its like i posted originally. I'm sure RadEn: i see what you are saying there, my trig is weak so i couldnt i denitrify that

OpenStudy (anonymous):

i meant identify

OpenStudy (raden):

ok, that problem can be : int (16*2sin(theta)*2cos(theta) d(theta) = 32 int (sin(2*theta)) d(theta)

OpenStudy (anonymous):

you lost me there, i was just going to find the antiderivative of 2cos(theta) since 16 is a constant

OpenStudy (raden):

= -16cos(2*theta) and then from x=2sin(theta), (i want to show in pic)